What's the differences between using the y=mx+b unlike the formula m=y2-y1/x2-x1 to find the slope?

Answer by nyc_function(2741)

A common form of a linear equation in the two variables x and y is y = mx + b,
where m and b are called constants. The origin of the name "linear" comes from the fact that the set of solutions of such an equation forms a straight line in the coordinate plane. In this particular equation, the constant m determines the slope or gradient of that line, and the constant term "b" determines the point at which the line crosses the y-axis, otherwise known as the y-intercept.
===============================================================================
In math, the slope or gradient of a line describes its steepness, incline, or grade. A higher slope value indicates a steeper incline.
Slope is normally described by the fraction of the "rise" divided by the "run" between two points on a given line or linear equation in the form y = mx + b.
The rise is represented by (y2 − y1).
The run is represented by (x2 − x1).
The slope m of the line is found using the formula m = (y2 - y1)/(x2 - x1).